OpenStudy (anonymous):

(12w^4-6w^3+w^2)/(3w^2 )

OpenStudy (anonymous):

\[{12w^4-6w^3+w^2 \over 3w^2}={12w^4 \over 3w^2}-{6w^3 \over 3w^2}+{w^2 \over 3w^2}\]Does this help you see how you can simplify it

OpenStudy (anonymous):

\[\huge {x^a \over x^b}=x^{a-b}\]

OpenStudy (jennychan12):

oh my. algebra comes back to haunt me... building off of @ChmE ... simplify each term... so you get \[4w^2 - 2w+\frac{ 1 }{ 3 }\] if you want to you can rewrite it as... \[\frac{ 12w^2-6w+1 }{ 3 }\] by multiplying each term by 3

OpenStudy (anonymous):

(12w^4)/(3w^2 ) - (6w^3)/(3w^2 ) + w^2/(3w^2 ) 4w^2 - how do I divide the 6w^3/3w^2

OpenStudy (anonymous):

I divide it leaves me with 1.5 on the exponent

OpenStudy (anonymous):

\[\huge {x^a \over x^b}=x^{a-b}\]

OpenStudy (anonymous):

\[\large {w^3 \over w^2}=w^{3-2}=w\]

OpenStudy (jennychan12):

in other words, you subtract the bottom exponent from the top exponent. you don't divide the exponents.

OpenStudy (anonymous):

okay I got it. Thanks